OpenStudy (anonymous):

Quadrilateral ABCD with vertices A(4, 3), B(4, -2), C(-4, -2) and D(-4, 3) is a rectangle, find the length of the diagonals.

OpenStudy (anonymous):

AB has length 5 AD has length 8 BD (diagonal) by Pythagorean Theorem has length \[\sqrt{8^{2}+5^{2}}=\sqrt{64+25}=\sqrt{89}\]